# envname-P envname#1

### v2

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) with external knowledge but still suffers from long contexts and disjoint retrieval–generation optimization. In this work, we propose CLaRa (Continuous Latent Reasoning), a unified framework that performs embedding-based compression and joint optimization in a shared continuous space. To obtain semantically rich and retrievable compressed vectors, we introduce SCP, a key-preserving data synthesis framework using QA and paraphrase supervision. CLaRa then trains the reranker and generator end-to-end via a single language modeling loss, with gradients flowing through both modules using a differentiable top-k estimator. Theoretically, this unified optimization aligns retrieval relevance with answer quality. Experiments across multiple QA benchmarks show that CLaRa achieves state-of-the-art compression and reranking performance, often surpassing text-based fine-tuned baselines.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful paradigm for enhancing large language models (LLMs) across diverse NLP tasks (Lewis et al., 2020; Gao et al., 2024; Li et al., 2024b; Wu et al., 2024; Abootorabi et al., 2025). By incorporating external evidence, RAG mitigates key weaknesses of LLMs such as hallucination (Ayala & Bechard, 2024) and knowledge obsolescence (Lau et al., 2025).
Most RAG systems suffer from a fundamental structural issue: retrieval and generation are optimized separately. Retrievers select documents based on surface-level similarity, while generators produce answers without providing feedback about what information is truly needed (Shi et al., 2025). This disjoint design leads to two intertwined challenges. (1) Optimization. Because document selection is discrete, gradients cannot flow from the generator back to the retriever (Sachan et al., 2021; Lin et al., 2024), hindering joint training and preventing the retriever from aligning with the generator’s task objective. (2) Efficiency. Dense retrievers rank documents in embedding space, whereas generators still consume raw text, resulting in an architectural mismatch. This mismatch yields (i) inconsistent representation spaces that prevent end-to-end optimization, (ii) redundant text processing that increases inference cost (Merth et al., 2024) and causes context overflow (Leng et al., 2024; Yue et al., 2025), and (iii) duplicated encoding for both retrieval and generation. Even if gradients could flow jointly, these inefficiencies would persist due to the lack of a shared latent space.
@@
Towards tackling this issue, we propose a unified framework that performs retrieval and generation over shared continuous document representations as shown in Fig 1. Instead of maintaining separate embeddings and raw text, we encode documents once into compact memory-token representations that serve both purposes. A central motivation behind this design is that supervised retrieval training typically relies on relevance-labeled data, which is scarce and often domain-specific. To overcome this limitation, we propagate the next-token prediction (NTP) loss from the generator to the retriever, providing a weakly supervised signal that naturally adapts retrieval to downstream generation objectives. This mechanism allows the retriever to learn which documents truly enhance answer generation rather than relying on surface-level similarity. Moreover, continuous representations and joint optimization are inherently complementary: continuous encodings make the retrieval process differentiable, while joint training aligns both modules within a shared semantic space optimized for reasoning.
This unified design resolves both challenges simultaneously. Optimization-wise, continuous representations enable differentiable top-$k$ selection via Straight-Through estimation, allowing generator gradients to update the retriever directly through gradient descent rather than inefficient RL sampling. Efficiency-wise, shared encodings remove redundant computation and drastically reduce context length, enabling fully end-to-end optimization and inference within the same representation space.
@@
To realize this vision, we present CLaRa (Continuous Latent Reasoning), a joint retrieval–generation framework built on shared compressed representations. In Stage I, we propose SCP (Salient Compressor Pretraining), which enhances semantic fidelity by constructing QA pairs that emphasize salient document content beyond surface reconstruction. In Stage II, CLaRa performs end-to-end joint training of the query encoder and answer generator under a unified next-token prediction loss, with differentiable top-$k$ selection via Straight-Through estimation. Theoretically, we show this unified objective yields valid gradients for retriever learning without explicit labels.
@@
We evaluate CLaRa on four single-hop and multi-hop QA benchmarks with Mistral-7B and Phi-4B. Results show that SCP produces semantically rich compressed representations, and CLaRa achieves state-of-the-art retrieval and generation performance—outperforming both supervised and unsupervised baselines, and even surpassing full-text fine-tuned models on several tasks.
@@
Figure 2: Overview of the SCP (Salient Compressor Pretraining) framework. It includes (a) synthetic data construction for pretraining, (2) compressor training using the pretraining data.
@@ 2 SCP: Salient Compressor Pretraining
Previous methods (Louis et al., 2025a; Cheng et al., 2025) typically use token-level reconstruction loss to learn doc representation. However, the learned representation may waste limited capacity/budget on token-by-token reconstruction which might be trivial. Also, the raw representation learned in such a way might not “digest” the document exhaustively. To enable the model to focus on abstracting and digesting semantically informative representations, we first synthesize pre-training data that highlights salient information. Based on this data, we train a compression framework, where a compressor learns to retain merely essential semantics (Figure 2).
@@ 2.1 Guided Data Synthesis for Semantic Preservation
We first construct a synthetic dataset where salient information is explicitly exposed through QA and paraphrasing. This way, the compressor later learns to identify, digest and retain the semantic core of the text by deeply processing the raw token-level information. As shown in Figure 2 (a), our synthetic data generation pipeline consists several steps: (1) salient information elicitation via QA and paraphrase generation, (2) automatic verification of coverage, and (3) regeneration of missing content.
@@
Using 2M sampled Wikipedia-2021 documents (Izacard et al., 2023), a locally deployed LLM (Qwen-32B) generates three complementary forms of supervision:
  • Simple QA: each pair captures a single fact, encouraging fine-grained factual retention. To avoid redundancy, the model is guided to extract distinct atomic facts that have not been covered by previous questions. This helps ensure broad coverage of salient information while keeping each question focused. Example: Q: “In which plant family is Trichocladus crinitus classified?” A: “Trichocladus crinitus belongs to the plant family Hamamelidaceae (the witch-hazel family).”
  • Complex QA: each pair integrates multiple facts, promoting relational reasoning and higher-level abstraction. During generation, the model is encouraged to connect previously unlinked facts, which reduces repetition and increases coverage of relational information. Example: Q: “Which player joined Newport County on 2 October 2014 and made his debut as a second-half substitute in a 1–0 defeat at Oxford United on 4 October?” A: “The player is James Loveridge.”
  • Paraphrase: paraphrased documents reorder sentences structure, thus altering the surface form of the text while preserving its core semantics. Learning such a mapping from original text to paraphrased text through an information bottleneck of continuous representation will enable the learned representation to focus on the semantics.
@@
Each document and its generated outputs (QA pairs or paraphrases) are verified by the same LLM for factual consistency and information coverage. When missing information is detected, the LLM reviews both the original text and existing QA pairs to generate additional ones capturing uncovered facts, iteratively up to ten rounds. Samples failing final coverage criteria are excluded. This iterative check ensures the model only learns from fully covered, factually faithful pairs.
@@ 2.2 Compressor Pretraining
Following PISCO (Louis et al., 2025a), we adopt a shared base model equipped with multiple LoRA adapters for modular control, where each adapter corresponds to a distinct function (compression or generation) as illustrated in Figure 2 (b).
@@
$M_{i}=\mathrm{LLM}_{\theta_{c}}([t_{1},\dots,t_{m},m_{1},\dots,m_{l}])[m{+}1:m{+}l].$
The compressed vector $M_{i}$ is concatenated with an instruction $I$ to form $T=[I;M_{i}]$. During pretraining, $I$ corresponds to general text-generation tasks (e.g., QA or paraphrasing), and later is replaced with task-specific prompts during instruction tuning. Only the generator LoRA $\theta_{g}$ is trained via cross-entropy loss:
$\mathcal{L}_{\text{CE}}(\theta_{c},\theta_{g})=-\sum_{(d_{i},I,R_{i}^{*})}\sum_{t=1}^{|R_{i}^{*}|}\log p_{\theta_{g}}\!\big(a_{i,t}^{*}\mid I,M_{i},a_{i,<t}^{*}\big).$ (2.1)
@@
The pretrained compressor is for general purpose. To adapt the compressor for downstream QA and also obtain an answer generator that can comprehend the continuous document representation, we optionally performed an additional instruction finetuning training. Note that this step is not necessary. It is for evaluation and ablation purpose, and we show the comparison in our experiments.
To achieve this, we use downstream training datasets in which the retrieved documents are paired with task instructions. These document–instruction pairs form the input to the model, while the output is a reference response generated by a teacher model conditioned on the same retrieved documents and instructions. Similar to the compressor pretraining stage, we jointly finetune the LoRA adapters of both the compressor and the answer generator during this instruction-tuning process.
@@ 3 CLaRa: Retrieval and generation joint training
As shown in Figure 3, each of the document is compressed into a dense embedding $\mathbf{M}_{i}=\theta_{c}(t_{i})$ using a pretrained compressor $\theta_{c}$. The compressor remains frozen to allow offline document encoding. We then train a query reasoner ($\theta_{qr}$), a LoRA adapter initialized from $\theta_{c}$, to represent queries in the same space and with the same number of memory tokens as document representations. Through next-token prediction (NTP) training, $\theta_{qr}$ learns not only to encode query intent but also to anticipate relevant document content, enhancing retrieval and answer generation. For example, given “Which city hosted the first modern Olympic Games?”, an embedding-based retriever may miss “first” or “modern,” whereas the NTP-trained query reasoner favors documents mentioning “Athens 1896,” which better satisfies retrieval with reasoning needs. We use cosine similarity to determine the relevance between query and documents:
@@
The top-$k$ document compressed embeddings $\{\mathbf{M}_{1},\dots,\mathbf{M}_{k}\}$ are concatenated with $Q$ and fed to the generator (activated by $\theta_{g}$), which produces the final answer. During training, both $\theta_{qr}$ and $\theta_{g}$ are updated via the unified language modeling loss:
$\mathcal{L}_{\text{CLaRa}}(\theta_{q},\theta_{g})=-\sum_{t=1}^{|R^{*}|}\log p_{\theta_{g}}\!\left(a_{t}^{*}\,\middle|\,Q,\mathbf{M}_{(1:k)},a_{<t}^{*}\right),$ (3.2)
@@
In the CLaRa framework, retrieval and generation are trained jointly, yet their connection is mediated by the top-$k$ selection of relevant documents. However, this discrete operation introduces a broken gradient problem: the generator’s supervision cannot propagate back to inform the retriever why certain documents should be preferred over others. To address this issue, we introduce top K selector via Straight-Through (ST) estimator, which conceptually acts as a “soft lens” — preserving the discrete retrieval behavior during inference while allowing smooth gradient feedback during training (see Algorithm 1 in Appendix for details).
Given cosine similarities $s=[s_{1},\dots,s_{D}]$, temperature $\tau$, and masking for previously selected items, the soft and hard selections are defined as:
@@
where $M\in\mathbb{R}^{B\times D\times d}$ is the matrix of all candidate embeddings with B denoting the batch size, D the number of candidate documents, and d the dimensionality of each document representation, typically defined as the product of the number of memory tokens and the hidden dimension of the underlying LLM.
@@
We provide a theoretical justification for why learning from NTP may yield stronger and more stable learning signals for the $\theta_{qr}$. Note that $\theta_{qr}$ can also be pre-trained independently using contrastive learning over positive and negative document–query pairs $(d^{+},q)$ and $(d^{-},q)$, following the DPR framework (Zhou & Chen, 2025). For query $x$ and document $d$, the retrieval score $s_{xd}$ defines
$p(d|x)=\frac{\exp(s_{xd}/\tau)}{\sum_{j}\exp(s_{xj}/\tau)},\quad p(y|x)=\sum_{d}p(d|x)p(y|x,d),\ \mathcal{L}=-\log p(y|x).$
When reranking and generation share same representations, $p(y|x,d)$ depends on $s_{xd}$, allowing generator gradients to flow into the reranker. With shared embeddings $r=\sum_{d}\pi_{d}(s)z_{d}$ and a Straight-Through estimator for top-$k$, the gradient is
$\frac{\partial\mathcal{L}}{\partial s_{xd}}=-\frac{1}{p(y|x)}\!\left[p(d|x)\!\big(p(y|x,d)-p(y|x)\big)+\frac{p(y|x,r)}{\tau}p(d|x)g^{\top}(z_{d}-r)\right],$
where $g=\nabla_{r}\log p(y|x,r)$.
Interpretation: Intuitively, the gradient coupling allows the retriever to receive two complementary learning signals. First, it is rewarded for ranking the correct documents higher through probabilistic alignment between $p(d|x)$ and $p(y|x,d)$. Second, it is guided to represent documents in a way that facilitates the generator’s reasoning via representation-level feedback from the gradient $g$. Together, these dual signals stabilize training—rather than oscillating between strong retrieval and weak generation, both modules progressively align within a shared semantic space, ensuring consistent retrieval–generation behavior (see Appendix A for details).
@@
To probe the information embedded within the query reasoner $\theta_{qr}$, we adopt the logit lens analysis technique (nostalgebraist, 2020). For each memory embedding, we project it through the LLM’s output head and record the top-50 tokens with the highest logits as topic tokens. We then aggregate and filter these decoded tokens to remove trivial elements such as punctuation or special symbols. As shown in Table 1, for the query “How many yards did the nephew of Ivory Lee Brown get during his 2004 true freshman season?”, the query embeddings decoded from the reasoner include the token “NFL”, “Oklahoma”, despite the fact that this word does not appear in the question itself. Interestingly, this token also occurs in the corresponding positive document and serves as a crucial clue for answering the question. This finding indicates that our end-to-end optimization enables the query reasoner to implicitly encode reasoning-relevant knowledge aligned with the gold evidence, thus enhancing retrieval accuracy and semantic alignment compared to baseline systems.
@@
Table 1: Analysis of Decoded Tokens from Query Reasoner via Logit Lens. The highlighted tokens (red) denote the new information reasoned by the query reasoner, while (blue) denotes key evidence for solving this multihop task.

### v3

@@ Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) with external knowledge but still suffers from long contexts and disjoint retrieval–generation optimization. In this work, we propose CLaRa (Continuous Latent Reasoning), a unified framework that performs embedding-based compression and joint optimization in a shared continuous space. To obtain semantically rich and retrievable compressed vectors, thereby reducing the document length fed into the generator, we introduce SCP, a key-preserving data synthesis framework based on question-answering and paraphrase supervision. CLaRa then trains the reranker and generator end-to-end via a single language modeling loss, with gradients flowing through both modules using a differentiable top-k estimator. Theoretically, this unified optimization aligns retrieval relevance with answer quality. Experiments across multiple QA benchmarks show that CLaRa achieves state-of-the-art compression and reranking performance, even at a text compression rate of 16, outperforming text-based fine-tuned baselines.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a powerful paradigm for enhancing LLMs across diverse NLP tasks (Lewis et al., 2020; Gao et al., 2024; Li et al., 2024b; Wu et al., 2024; Abootorabi et al., 2025). By incorporating external evidence, RAG mitigates key weaknesses of LLMs such as hallucination (Ayala & Bechard, 2024) and knowledge obsolescence (Lau et al., 2025).
Most RAG systems suffer from a fundamental structural issue: retrieval and generation are optimized separately. Retrievers select documents based on surface-level similarity, while generators produce answers without providing feedback about what information is truly needed (Shi et al., 2025). This leads to two intertwined challenges. (1) Efficiency. Dense retrievers rank documents in embedding space, while generators still consume raw text, resulting in an architectural mismatch. This mismatch yields (i) inconsistent representation spaces that prevent end-to-end optimization, (ii) redundant text processing that increases inference cost (Merth et al., 2024) and causes context overflow (Leng et al., 2024; Yue et al., 2025), and (iii) duplicated encoding for both retrieval and generation. Even if gradients could flow jointly, these inefficiencies would persist due to the lack of a shared latent space. (2) Optimization. Because document selection is discrete, gradients cannot flow from the generator back to the retriever (Sachan et al., 2021; Lin et al., 2024), hindering joint training and preventing the retriever from aligning with the generator’s task objective.
@@
To address the aforementioned issues, we propose a unified framework that performs retrieval and generation over shared continuous document representations, as illustrated in Fig. 1. Each document is encoded only once into a compact set of memory tokens, which simultaneously serve both retrieval and generation. This design is motivated by: (1) Not all tokens in textual documents carry informative content. Text can be compressed into continuous representations with high representational capacity that preserve only the most salient information, and more compact document representations can also effectively reduce the input length of the generator. (2) Moreover, continuous representations and joint optimization are inherently complementary: continuous encodings make the retrieval process differentiable, while joint training aligns the retriever and the generator within a shared semantic space optimized for reasoning.
We design a salient-information–aware pretraining objective that pushes the model to compress documents into compact encoded representations. The generator can then answer questions by relying solely on these shorter input representations. Meanwhile, we backpropagate the generator’s next-token prediction (NTP) loss to the retriever, providing a weak supervision signal that naturally adapts retrieval to downstream generation objectives. This mechanism enables the retriever to learn which documents truly contribute to answer generation, rather than relying only on surface-level similarity. This unified design simultaneously addresses both challenges. Efficiency-wise, shared encodings eliminate redundant computation, making true end-to-end optimization and inference within a unified representation space possible. Moreover, during generation, the use of shorter context lengths saves precious resources when large language models process long-form inputs. Optimization-wise, continuous representations enable differentiable top-$k$ selection via Straight-Through (ST) estimation (Bengio et al., 2013), allowing generator gradients to update the retriever directly through gradient descent rather than inefficient RL sampling.
@@
To realize this vision, we present CLaRa (Continuous Latent Reasoning), a joint retrieval–generation framework built on shared compressed representations. In Stage I, we propose SCP (Salient Compressor Pretraining), which enhances semantic fidelity by constructing QA pairs that emphasize salient document content beyond surface reconstruction. In Stage II, CLaRa performs end-to-end joint training of the query encoder and answer generator under a unified next-token prediction loss, with differentiable top-$k$ selection via ST estimation. Theoretically, we show this unified objective yields valid gradients for retriever learning without explicit labels.
@@
We evaluate CLaRa on four single-hop and multi-hop QA benchmarks with Mistral-7B and Phi-4B. Results show that SCP produces semantically rich compressed representations, and CLaRa achieves state-of-the-art retrieval and generation performance—outperforming both supervised and unsupervised baselines, and even surpassesing text-only DRO methods when our text compression ratio is 16.
@@
Figure 2: Overview of the SCP framework, featuring (a) synthetic pretraining data construction; (b) compressor training using this data.
@@ 2 SCP: Salient Compressor Pretraining
Previous methods (Louis et al., 2025a; Cheng et al., 2025) typically use token-level reconstruction loss to learn document representation. This risks wasting limited capacity/budget on potentially trivial token-by-token reconstruction, and the resulting representations might not “digest” the document in a way that retains all relevant information. To enable the model to focus on generating semantically informative representations, we first synthesize pre-training data that highlights salient information. Based on this data, we train a compression framework, where a compressor learns to retain just the essential semantics (Fig 2).
@@ 2.1 Guided Data Synthesis for Semantic Preservation
We first construct a synthetic dataset where salient information is explicitly exposed through QA and paraphrasing. This way, the compressor later learns to identify, digest and retain the semantic core of the text by deeply processing the raw token-level information. As shown in Fig 2 (a), our synthetic data generation pipeline consists several steps: (1) salient information elicitation via QA and paraphrase generation, (2) automatic verification of coverage, and (3) regeneration of missing content.
@@
Using 2M sampled Wikipedia-2021 documents (Izacard et al., 2023), a locally deployed LLM (Qwen-32B) generates three complementary supervision signals:
  • Simple QA: each pair targets a single atomic fact to encourage fine-grained factual retention, while explicitly avoiding redundancy by extracting distinct facts not covered by previous questions.
  • Complex QA: each pair integrates multiple facts to promote relational reasoning, with generation guided to connect previously unlinked information and increase coverage.
  • Paraphrase: paraphrased documents alter surface structure while preserving semantics, encouraging the model to learn semantic representations through a continuous information bottleneck.
@@
Each document and its generated outputs (QA pairs or paraphrases) are verified by the locally deployed LLM for factual consistency and information coverage. When missing information is detected, the LLM reviews both the original text and existing QA pairs to generate additional ones capturing uncovered facts, iteratively up to ten rounds. Samples failing final coverage criteria are excluded. This iterative check ensures the model only learns from fully covered, factually faithful pairs.
@@ 2.2 Compressor Pretraining
Following PISCO (Louis et al., 2025a), we adopt a shared base model equipped with multiple LoRA adapters for modular control, where each adapter corresponds to a distinct function (compression or generation) as shown in Fig 2 (b).
@@
$M_{i}=\mathrm{LLM}_{\theta_{c}}([t_{1},\dots,t_{m},m_{1},\dots,m_{l}])[m{+}1:m{+}l],$
which is concatenated with an instruction $I$ to form $T=[I;M_{i}]$. During pretraining, $I$ corresponds to general text-generation tasks (e.g., QA or paraphrasing); during instruction tuning it is replaced with task-specific prompts. Only the generator LoRA $\theta_{g}$ is trained via cross-entropy loss:
$\mathcal{L}_{\text{CE}}(\theta_{c},\theta_{g})=-\sum_{(d_{i},I,R_{i}^{*})}\sum_{t=1}^{|R_{i}^{*}|}\log p_{\theta_{g}}\!\big(a_{i,t}^{*}\mid I,M_{i},a_{i,<t}^{*}\big).$
@@
To adapt the general-purpose pretrained compressor for downstream QA, and also obtain an answer generator that can comprehend the continuous document representation, we optionally performed an additional instruction finetuning training. To achieve this, we use downstream training datasets in which the retrieved documents are paired with task instructions. These document–instruction pairs form the input to the model, while the output is a reference response generated by a teacher model conditioned on the same retrieved documents and instructions. Similar to the compressor pretraining stage, we jointly finetune the LoRA adapters of both the compressor and the answer generator during this instruction-tuning process.
@@ 3 CLaRa: Retrieval and generation joint training
As shown in Fig 3, each document is compressed into a dense embedding $\mathbf{M}_{i}=\theta_{c}(t_{i})$ using the pretrained compressor $\theta_{c}$. The compressor remains frozen to allow offline document encoding. We then train a query reasoner ($\theta_{qr}$), a LoRA adapter initialized from $\theta_{c}$, to represent queries in the same space and with the same number of memory tokens as document representations. Through NTP training, $\theta_{qr}$ learns not only to encode query intent but also to anticipate relevant document content, enhancing retrieval and answer generation. For example, given “Which city hosted the first modern Olympic Games?”, an embedding-based retriever may miss “first” or “modern,” whereas the NTP-trained query reasoner favors documents mentioning “Athens 1896,” which better satisfies retrieval with reasoning needs. We use cosine similarity between query embedding $\mathbf{q}_{i}$ and document embedding $\mathbf{M}_{i}$ to obtain a relevance score:
@@
The top-$k$ scoring embeddings $\{\mathbf{M}_{1},\dots,\mathbf{M}_{k}\}$ are concatenated with $\mathbf{q}$ and fed to the generator ($\theta_{g}$), which produces the final answer. During training, both $\theta_{qr}$ and $\theta_{g}$ are updated via the unified language modeling loss:
$\mathcal{L}_{\text{CLaRa}}(\theta_{q},\theta_{g})=-\sum_{t=1}^{|R^{*}|}\log p_{\theta_{g}}\!\left(a_{t}^{*}\,\middle|\,\mathbf{q},\mathbf{M}_{(1:k)},a_{<t}^{*}\right),$ (3.2)
@@
In the CLaRa framework, retrieval and generation are trained jointly, yet their connection is mediated by the top-$k$ selection of relevant documents. However, this discrete operation introduces a broken gradient problem: the generator’s supervision cannot propagate back to inform the retriever why certain documents should be preferred over others.
Previous methods such as DDR-RAG (Li et al., 2024a) update the retriever by performing multiple retrieval samplings and using rewards derived from the generator’s outputs under different document samples. However, this RL-based approach suffers from training instability and low efficiency due to repeated sampling. Instead, we use ST estimation for top-$k$ selection, which conceptually acts as a “soft lens” — preserving the discrete retrieval behavior during inference while allowing smooth gradient feedback during training (see Algorithm 1 in Appendix for details).
Given cosine similarities $s_{b}=[s_{1},\dots,s_{D}]$ for the $b$th document set in a batch, temperature $\tau$, and masking for previously selected items, the soft and hard selections are defined as:
@@
where $M\in\mathbb{R}^{B\times D\times E}$ is the matrix of all candidate embeddings with B denoting the batch size, D denoting the number of candidate documents, and E denoting the dimensionality of each document representation, typically defined as the product of the number of memory tokens and the hidden dimension of the underlying LLM.
@@
We provide an explanation for why learning from NTP yields stronger and more stable training signals for the retriever. By coupling retrieval and generation through shared representations and ST-based top-$k$ selection, the retriever receives two complementary learning signals. First, it is encouraged to rank documents that better support correct generation, aligning retrieval probabilities with generation outcomes. Second, it receives representation-level feedback from the generator, guiding document embeddings toward being more useful for downstream reasoning. This dual feedback stabilizes joint training by progressively aligning retrieval and generation within a shared semantic space (see Appendix A for details).
@@
To probe the information embedded within the query reasoner $\theta_{qr}$, we adopt the logit lens analysis technique (nostalgebraist, 2020). For each memory embedding, we project it through the LLM’s output head and record the top-50 tokens with the highest logits as topic tokens. We then aggregate and filter these decoded tokens to remove trivial elements such as punctuation or special symbols. As shown in Fig. 4, for the query “How many yards did the nephew of Ivory Lee Brown get during his 2004 true freshman season?”, the query embeddings decoded from the reasoner includes the tokens “NFL”, “Oklahoma”, despite the fact that these word do not appear in the question itself. Interestingly, these token do occur in the corresponding positive document and serve as a crucial clue for answering the question. This finding indicates that our end-to-end optimization enables the query reasoner to implicitly encode reasoning-relevant knowledge aligned with the gold evidence, thus enhancing retrieval accuracy and semantic alignment compared to baseline systems.
@@
Figure 4: Analysis of Decoded Tokens from Query Reasoner via Logit Lens. The highlighted tokens (red) denote the new information reasoned by the query reasoner, while (blue) denotes key evidence for solving this multihop task.